# CliniBench: A Clinical Outcome Prediction Benchmark for Generative and Encoder-Based Language Models

### v1

@@ 5 Results
Zero-Shot Zero-Shot + CoT Few-Shot 5 Demonstrations
Data Model Rec. Prec. MAP MD Acc. Rec. Prec. MAP MD Acc. Rec. Prec. MAP MD Acc.
hosp
Encoder-based Classifier
BioMedBERT-110M 30.74 19.91 33.61 78.21 -
BioMedBERT-110M-tuned 22.91 29.55 19.34 59.16
GatorTronS-345M 30.45 22.70 35.26 80.94
GatorTronS-345M-tuned 28.61 33.44 22.06 65.22
M2-80M 14.90 08.51 24.34 60.83
M2-80M-tuned 11.72 12.05 14.80 40.47
@@
Llama 3.1-8B 16.61 12.16 12.88 35.22 14.85 10.89 11.49 37.96 21.76 16.00 21.33 47.23
Llama 3.1-8B-Instruct 22.31 10.88 14.05 47.53 20.84 09.93 11.95 49.61 20.99 14.24 18.06 44.34
Llama 3.1-70B 17.29 09.78 11.40 35.21 20.72 09.42 12.36 43.63 20.55 12.67 19.68 49.05
Llama 3.1-70B-Instruct 23.04 10.40 12.48 52.20 23.70 10.00 12.53 51.39 21.89 14.43 18.02 45.49
Mistral-7B 10.45 12.02 10.12 27.25 00.40 03.86 02.29 00.69 19.15 13.78 18.02 44.13
Mistral-7B-Instruct 26.45 10.52 14.88 54.06 23.96 09.46 12.81 49.76 25.33 14.29 18.77 48.89
Mistral-Nemo-13B 14.87 11.30 12.53 32.83 13.66 09.37 09.90 38.32 19.18 13.76 18.44 45.97
Mistral-Nemo-13B-Instruct 22.23 09.55 12.34 50.35 21.19 09.22 10.44 51.61 21.15 15.28 19.09 48.41
Qwen2-7B 22.86 11.32 15.12 43.37 21.04 11.62 14.92 42.01 17.70 14.94 18.81 43.40
Qwen2-7B-Instruct 25.80 11.27 15.28 48.95 24.63 10.76 15.32 49.38 24.32 13.68 19.38 51.65
Qwen2-72B 24.63 11.06 15.45 46.17 22.18 10.42 14.86 44.05 22.57 14.36 20.38 49.44
Qwen2-72B-Instruct 27.24 11.09 17.28 51.58 25.66 10.91 15.28 50.24 25.29 15.82 21.51 53.00
icu
Encoder-based Classifier
BioMedBERT-110M 19.60 17.99 32.44 72.83 -
BioMedBERT-110M-tuned 15.56 23.13 20.03 55.91
GatorTronS-345M 28.61 23.96 35.66 78.44
GatorTronS-345M-tuned 20.00 29.62 21.13 63.66
M2-80M 10.16 08.27 23.12 56.05
M2-80M-tuned 08.13 08.98 10.72 35.05
Generative LLM
Llama 3.1-8B 13.97 14.81 13.50 33.43 12.25 13.41 12.36 35.57 14.12 14.60 17.62 48.33
Llama 3.1-8B-Instruct 19.59 13.93 15.63 47.34 17.88 13.00 13.21 47.65 15.52 14.36 15.93 44.27
Llama 3.1-70B 15.40 12.44 12.30 33.60 18.73 11.87 12.98 40.81 13.74 12.10 16.54 50.85
Llama 3.1-70B-Instruct 20.31 13.36 14.70 50.37 21.24 12.66 15.85 50.18 17.06 15.38 16.24 44.90
Mistral-7B 08.84 13.49 09.94 25.66 00.34 02.63 04.58 01.38 11.08 11.92 14.40 43.16
Mistral-7B-Instruct 23.65 13.67 16.86 51.37 21.59 12.12 14.64 46.10 20.96 15.08 16.98 47.72
Mistral-Nemo-13B 13.49 14.31 13.55 31.63 11.11 11.63 10.86 37.51 11.31 12.73 15.13 44.95
Mistral-Nemo-13B-Instruct 19.64 11.97 13.87 48.39 19.03 11.96 12.46 48.54 11.24 13.35 15.24 43.28
Qwen2-7B 20.95 13.92 15.54 40.97 19.05 13.96 15.49 40.19 11.23 13.36 15.80 43.35
Qwen2-7B-Instruct 23.06 14.14 15.33 45.86 22.13 13.89 15.18 46.30 18.23 14.40 17.61 50.52
Qwen2-72B 22.55 14.06 15.89 43.28 20.32 13.14 15.57 40.93 15.27 13.80 17.39 49.40
Qwen2-72B-Instruct 24.11 14.20 16.82 48.11 22.78 14.10 16.47 47.16 17.42 15.74 18.75 52.06
Table 2: (MAP: Mean Average Precision, MD Acc.: Main Diagnosis Accuracy) Zero-shot performance on the hosp and icu splits over the top-20 predictions. The reported metrics are macro averaged. We aggregate ICD-9 and ICD-10 and report an unweighted mean. The dashed lines separate the encoder from the generative LLMs. CoT is not beneficial while instruct fine-tuning boosts performance. Threshold tuning decreases the performance of encoder models within the top-20 predictions. We depict the best performing models in bold and underline the best performing generative models in each category.
Table˜2 highlights key differences between encoder-based and generative LLM architectures on the hosp and icu splits.
@@ 5.2 Retrieval Models
Table˜8 in the appendix shows the performance of retrieval models using majority voting on the diagnosis prediction task. We observe that our fine-tuned classifier, BioMedBERTft, achieves the best performance. Notably, the baseline BM25 outperforms all retrieval models based on semantic similarity. Among the sentence-transformer models, performance differences are minimal, with PubMedBERTNeuML outperforming others. These findings suggest that using domain specific retrieval models provides a modest advantage for clinical tasks.
@@
hosp Zero-Shot 0 27.24 11.09 17.28 51.58
@@
icu Zero-Shot 0 24.11 14.20 16.82 48.11
Random 1 17.67 13.95 16.06 40.80
3 15.90 13.27 16.00 41.22
5 14.24 13.45 15.77 40.66
BM25 1 19.97 15.29 17.46 47.27
3 18.89 15.89 19.09 51.18
5 17.42 15.73 20.48 52.06
PubMedBERTNeuML 1 18.25 13.83 17.16 45.59
3 17.13 13.97 18.43 50.50
5 16.10 14.48 19.92 52.34
BioMedBERTft 1 18.97 14.60 17.69 44.80
3 18.04 14.86 18.96 50.92
5 18.12 15.44 19.57 52.59
Gold Heuristic 1 21.31 17.79 19.75 48.45
3 22.72 20.24 26.88 59.15
5 22.83 21.21 29.33 61.66
@@ 5.3 Few-Shot
As shown in Fig.˜2 and Table˜2, providing generative LLMs with even a single example enhances their performance compared to zero-shot settings. This improvement is likely because the demonstration helps align the model more closely with the task, particularly benefiting non-instruct variants of generative models. Adding more demonstrations continues to boost performance, albeit with diminishing gains.
@@
The gold heuristic provides an idealized upper boundary for performance by retrieving the most relevant samples (see Table˜8). It achieves the highest metrics across both splits, with the main diagnosis accuracy reaching 62.62% for hosp and 61.66% for icu, along with substantial MAP improvements. Among the retrievers, BM25 emerges as a robust baseline, consistently delivering competitive MAP and main diagnosis accuracy, outperforming NeuML-PubMedBERT and BioMedBERTft in several configurations. Task-specific fine-tuning provides a slight advantage for BioMedBERTft, which outperforms NeuML-PubMedBERT. Random retrieval degrades the performance of generative models, underscoring the importance of relevance-driven retrieval strategies for downstream tasks.
@@
Figure 3: Macro recall and precision by model and classes grouped into tertiles by class frequency aggregated over ICD-9 and ICD-10 splits. Diagram shows the icu (top) and the hosp (bottom) splits.
The results in Fig.˜3 highlight differences in performance across labels with different frequencies (see Table˜1). We observe that general performance is better for frequent labels. We find that retrieval augmentation boosts precision and reduces recall for all tertiles, particularly for less frequent labels. Generative models achieve a higher recall for rare codes than encoder-based classifiers. Finally, we observe that threshold tuning GatorTronS improves recall in the tail, while decreasing it in the head tertile.
@@
We observe that the number of parameters does not have a strong impact on performance. Further, we find that few-shot learning especially benefits smaller models, increasing the performance drastically in contrast to larger models. We hypothesize that guided decoding reduces the necessity of capacity in the larger models, making small models suitable for the task. Furthermore, we find that instruct fine-tuning has a larger effect on performance compared to the number of parameters (see Fig.˜10), narrowing the gap between small and large models.
@@ 5.4 Error Analysis
In addition to presenting quantitative results, we conduct a qualitative analysis of 20 randomly sampled examples with an especially low macro recall per experiment from the Qwen2-72B and Qwen2-72B-Instruct in zero-shot, few-shot (BM-25), and chain-of-thought settings. From our analysis, we deduce the following error classes:
@@
We observe that generative models sometimes tend to repeat certain outputs or generate text with little to no variance (e.g. Lung Cancer, Lung Carcinoma, Cancer…). This leads to a lower number of predicted codes after mapping and deduplication. The results indicate that this behavior mainly appears for few-shot augmentation as 80% of conducted one-shot predictions and 75% of five-shot prediction have a low variance output. The results reported in Fig.˜4 suggest that this behavior exists for all models, including the zero-shot scenario, because all models predict less than the expected 20 codes on average. Furthermore, we observe that chain-of-thought leads to similar results, with 50% of the examples that fall under this error class. Finally, we find that non-instruct fine-tuned models generate more varied output and thus more diverse diagnoses.
@@
Furthermore, we analyzed how often the generative LLMs produce valid but irrelevant text (e.g. procedures instead of diagnoses) or non-meaningful text at all (diagnosis 1, diagnosis 2, diagnosis 3 …). We observe that especially in the zero-shot case and for non-instruct fine-tuned models 65% of the errors originate from unrelated but still meaningful output. In the analyzed examples, we often find that the model outputs procedures instead of diagnoses or copies parts of the shown admission note. We find that a single few-shot example specifically increases the number of non meaningful generated text in 20% of the cases. This could imply that the model focuses on structural elements of the provided demonstration instead of task-specific information.
Text-to-class mapping.
We find that generative models often output diagnostic descriptions that exactly match the short or long descriptions of the respective ICD codes. This allows us to independently compare the performance of exact matched descriptions and codes that are mapped via the class mapper. For Qwen2-72B Instruct in the zeroshot scenario, only 5% of generated descriptions match ICD codes exactly; with 1, 3, and 5 few shots, this number rises to 32%, 51%, and 63% respectively. For the non-instruct fine-tuned version of the model, the number of exact matches rises even further to 56%, 70% and 76%. This indicates an adaption to the task through the provided few-shot examples. Furthermore, it shows that the model has seen the respective ICD short or long descriptions during pre-training and can correctly replicate them when it is provided with enough context. Among 5-shot predictions of Qwen2-72B Instruct, directly matched codes are correct 41% of the time, while mapped ones reach 20%. For Qwen2-7B, mapped codes retain 94% of the accuracy of directly matched ones (30% vs. 32%). While the accuracy of mapped codes is lower, their generated text contains enough relevant information that enables the class-mapper to resolve the correct code. This gap in performance is closely tied to the quality of the generated text. Accurate mappings result from meaningful clinical descriptions (e.g., Hypertensive heart disease, Hypotension), while nonsensical outputs (e.g., diagnosis 1, diagnosis 2, ___ etc.) almost never yield correct codes.
Performance drop for longer inputs.
Figure 5: Micro $F_{1}$ over different sequence lengths averaged over all dataset splits. Encoder models have a drastically reduced performance with sequences >850 tokens while the generative Models tend have a more consistent performance on longer sequences. The token count varies across models due to differences in tokenizers.
6 Discussion and Future Work
Our experiments demonstrate that the fine-tuned encoder-based models outperform current open source generative LLMs for diagnosis prediction, even when supported with strategies such as RAG or CoT prompting.
6.1 Methodological Improvements
We observe two main problems for the generative models: First, the output lacks diversity and second, retrieval augmentation does not reach its full potential. To address these problems, we suggest improvements in the following three key areas: output verification, more effective retrieval strategies, and different sampling strategies.
Output verification for meaningful outputs.
We observe that generative models, especially in the few-shot context, generate outputs with low variance or unrelated output. To address this issue, we suggest adding an LLM based verification mechanism, such as the one proposed by Gero et al. (2023). This mechanism could help to find missed predictions, filter out unrelated ones, or reduce the number of duplicate predictions.
Probabilistic sampling for diverse outputs.
Another option to increase the variance in the generated output is to use other sampling strategies than greedy sampling. Alternatives such as top-p or top-k sampling, as well as beam search, could help to generate more diverse predictions. However, probabilistic sampling requires multiple iterations with different seeds to obtain statistically significant improvements, which incurs high computational costs.
Beyond query-by-document retrieval.
In our experiments, latent outcome similarity outperforms semantic similarity for retrieval augmentation (see Table˜3). The gold heuristic surpasses both but uses ground-truth labels instead of the text of admission notes for retrieval. This indicates that more information than the text is required to retrieve relevant information to augment the generative models. Therefore, it could be beneficial to augment the query and knowledge base with other available internal clinical information such as laboratory results or imaging data.
Domain adaptation.
From the results of our CoT experiments we deduce that there is a lack of data that reflects clinical reasoning. To address this limitation, we suggest adapting the model using clinical guidelines. Furthermore, we encourage the creation of custom datasets, similar to the one proposed by Wang et al. (2025), that outline clinicians’ thought processes during differential diagnosis. In consequence, the integration of this data could also improve the effectiveness of CoT due to improved reasoning.
Challenges in model evaluation.
For the task of diagnosis prediction from admission notes, previous work primarily reported threshold-free metrics, such as AUROC or PR-AUC, which often yield high scores. However, diagnosis prediction in real-world applications would benefit from discretized or ranked predictions. To our knowledge, we are the first to report threshold-based metrics for this task and find that the scores are generally low (see Table˜2). This indicates the necessity for substantial methodological improvements to enhance model reliability and clinical applicability.
6.2 Cost Analysis
Selecting the most cost-effective model architecture depends heavily on the specific application scenario. In many medical contexts, models are required that can continually be fine-tuned to rapidly adapt to new diseases or treatments. For example, fine-tuning a BERT-base-sized model on approximately 330,000 documents requires approx. 0.3 PFLOPs Kaplan et al. (2020) per epoch. Fine-tuning in our experiments requires at most 20 epochs, amounting to a total of 8.4 PFLOPs. This computational budget would allow to evaluate around 20% of the whole dataset with a 70B model (1 PFLOPs). An 8B generative model needs 9.2 PFLOPs for the same evaluation. Fine-tuning the BERT-base model and evaluating it on the test set has similar costs as evaluating the test set with a 70B model. The inference cost of BERT is in comparison negligible (0.2 TFLOPs/example). However, this estimate ignores several real-world considerations, such as downstream task performance and infrastructure requirements. Larger generative models demand substantially more hardware during inference, which might remain underutilized between tasks. Nevertheless, these costs could be partially offset by leveraging techniques such as pruning and quantization.
6.3 Ethical Considerations
The use of language models for diagnosis prediction raises important ethical concerns. For example, inaccurate or biased predictions can negatively impact patient care, as admission notes may lack critical information. It is important to note that an automatic diagnosis prediction system should never be seen as a replacement for medical personnel. Rather, it should be used as a supporting tool in the differential diagnosis process. Another major ethical concern is data privacy, as third parties data access can lead to serious consequences, including financial harm, discrimination, and compromised care. Finally, there is a risk that ICD code prediction might be misused to maximize profits, prioritizing financial incentives over patient care. Addressing these concerns is vital for the safe and effective use of AI in healthcare in the future.

### v2

@@ 5 Results
Zero-Shot prompted Zero-Shot prompted + CoT Few-Shot 5 Demonstrations
Data Model Rec. Prec. MAP MD Acc. F1 Rec. Prec. MAP MD Acc. F1 Rec. Prec. MAP MD Acc. F1
HOSP Encoder-based Models
BioMedBERT-110M 30.74 19.91 33.61 78.21 14.63
BioMedBERT-110M-tuned 22.91 29.55 19.34 59.16 21.74
GatorTronS-345M 30.45 22.70 35.26 80.94 18.46
GatorTronS-345M-tuned 28.61 33.44 22.06 65.22 25.96
M2-80M 14.90 08.51 24.34 60.83 4.54
M2-80M-tuned 11.72 12.05 14.80 40.47 11.16
@@
Llama 3.1-8B 16.61 12.16 12.88 35.22 11.06 14.85 10.89 11.49 37.96 09.68 21.76 16.00 21.33 47.23 12.41
Llama 3.1-8B-Instruct 22.31 17.23 14.05 47.53 11.61 20.84 09.93 11.95 49.61 10.11 20.99 14.24 18.06 44.34 15.45
Llama 3.1-70B 17.29 09.78 11.40 35.21 09.66 20.72 09.42 12.36 43.63 10.20 20.55 12.67 19.68 49.05 15.08
Llama 3.1-70B-Instruct 23.04 10.40 12.48 52.20 11.32 23.70 10.00 12.53 51.39 10.59 21.89 14.43 18.02 45.49 15.82
Mistral-7B 10.45 12.02 10.12 27.25 08.50 00.40 03.86 02.29 00.69 00.15 19.15 13.78 18.02 44.13 15.29
Mistral-7B-Instruct 26.45 10.52 14.88 54.06 12.16 23.96 09.46 12.81 49.76 10.86 25.33 14.29 18.77 48.89 15.93
Mistral-Nemo-13B 14.87 11.30 12.53 32.83 10.15 13.66 09.37 09.90 38.32 08.00 19.18 13.76 18.44 45.97 15.10
Mistral-Nemo-13B-Instruct 22.23 09.55 12.34 50.35 10.38 21.19 09.22 10.44 51.61 09.57 21.15 15.28 19.09 48.41 16.05
Qwen2-7B 22.86 11.32 15.12 43.37 12.26 21.04 11.62 14.92 42.01 12.04 17.70 14.94 18.81 43.40 15.31
Qwen2-7B-Instruct 25.80 11.27 15.28 48.95 12.90 24.63 10.76 15.32 49.38 12.15 24.32 13.68 19.38 51.65 15.54
Qwen2-72B 24.63 11.06 15.45 46.17 12.47 22.18 10.42 14.86 44.05 11.49 22.57 14.36 20.38 49.44 16.46
Qwen2-72B-Instruct 27.24 11.09 17.28 51.58 13.13 25.66 10.91 15.28 50.24 12.75 25.29 15.82 21.51 53.00 17.57
Table 2: (MAP: Mean Average Precision, MD Acc.: Main Diagnosis Accuracy) Zero-shot performance on the HOSP split over the top-20 predictions. The reported metrics are macro averaged. We aggregate ICD-9 and ICD-10 and report an unweighted mean. CoT is not beneficial while instruct fine-tuning boosts performance. Threshold tuning decreases the performance of encoder models within the top-20 predictions. We depict the best performing models in bold and underline the best performing generative models in each category.
@@ 5.2 Retrieval Models
Table˜5 in the appendix shows the performance of retrieval models using majority voting on the diagnosis prediction task. We observe that our fine-tuned classifier, BioMedBERTft, achieves the best performance. Notably, the baseline BM25 outperforms all retrieval models based on semantic similarity. Among the sentence-transformer models, performance differences are minimal, with PubMedBERTNeuML outperforming others. These findings suggest that using domain specific retrieval models provides a modest advantage for clinical tasks.
@@
HOSP Zero-Shot prompted 0 27.24 11.09 17.28 51.58
@@ 5.3 Few-Shot
As shown in Table˜2, Table˜3 and Fig.˜4 (ICU scores reported in Appendix Table˜11 and Table˜10), providing generative LLMs with even a single example enhances their performance compared to zero-shot settings. This improvement is likely because the demonstration helps align the model more closely with the task, particularly benefiting non-instruct variants of generative models. Adding more demonstrations continues to boost performance, albeit with diminishing gains.
@@
The gold heuristic provides an idealized upper boundary for performance by retrieving the most relevant samples (see Table˜5). It achieves the highest metrics across both splits, with the main diagnosis accuracy reaching 62.62% for HOSP and 61.66% for ICU, along with substantial MAP improvements. Among the retrievers, BM25 emerges as a robust baseline, consistently delivering competitive MAP and main diagnosis accuracy, outperforming NeuML-PubMedBERT and BioMedBERTft in several configurations. Task-specific fine-tuning provides a slight advantage for BioMedBERTft, which outperforms NeuML-PubMedBERT. Random retrieval degrades the performance of generative models, underscoring the importance of relevance-driven retrieval strategies for downstream tasks.
@@
Figure 2: Macro recall and precision by model and classes grouped into tertiles by class frequency aggregated over both HOSP splits (ICD-9 and ICD-10).
The results in Fig.˜2 (see also Appendix Fig.˜6) highlight variations in performance across labels with different frequencies. We observe that general performance is better for frequent labels. We find that retrieval augmentation boosts precision and reduces recall for all tertiles, particularly for less frequent labels. Generative models achieve a higher recall for rare codes than encoder-based classifiers. Finally, we observe that threshold tuning GatorTronS improves recall in the tail, while decreasing it in the head tertile.
@@
We observe that the number of parameters does not have a strong impact on performance. Further, we find that few-shot learning especially benefits smaller models, increasing the performance drastically in contrast to larger models. We hypothesize that guided decoding reduces the necessity of capacity in the larger models, making small models suitable for the task. Furthermore, we find that instruct fine-tuning has a larger effect on performance compared to the number of parameters (see Fig.˜9), narrowing the gap between small and large models.
@@ 5.4 Error Analysis
In addition to presenting quantitative results, we conduct a qualitative analysis of 20 randomly sampled examples with especially low macro recall from Qwen2-72B and Qwen2-72B-Instruct in zero-shot, few-shot (BM-25), and chain-of-thought settings. From our analysis, we deduce the following error classes:
@@
We observe that generative models sometimes generate text with little to no variance (e.g., Lung Cancer, Lung Carcinoma, Cancer), leading to fewer predicted codes after deduplication. This behavior appears mainly in few-shot augmentation (80% of one-shot, 75% of five-shot predictions) and 50% of CoT predictions. As shown in Fig.˜3, all models predict fewer than the expected 20 codes on average, suggesting this issue exists across scenarios. Non-instruct fine-tuned models generate more varied output and thus more diverse diagnoses.
@@
We find that LLMs frequently produce valid but irrelevant text (e.g., procedures instead of diagnoses) or non-meaningful text (e.g., diagnosis 1, diagnosis 2, diagnosis 3). In zero-shot settings with non-instruct models, 65% of errors originate from unrelated meaningful output, where models often output procedures or copy parts of the admission note. Interestingly, a single few-shot example increases non-meaningful text generation in 20% of cases, suggesting the model focuses on structural rather than task-specific elements.
6 Future Work
To address the limitations identified in our study, we propose several directions for future research that could substantially improve the performance of generative models for clinical outcome prediction.
Output Verification and Sampling
To address low-variance and unrelated outputs, we suggest adding LLM-based verification mechanisms Gero et al. (2023) to filter predictions and reduce duplicates. Additionally, probabilistic sampling strategies (top-p, top-k, beam search) could increase output diversity.
Enhanced Retrieval Strategies
Given that latent outcome similarity outperforms semantic similarity and the gold heuristic surpasses both, future work should augment retrieval with additional clinical information beyond admission notes, such as laboratory results or imaging data, to improve demonstration selection for few-shot learning.
Domain Adaptation and Continual Learning
To enable effective chain-of-thought reasoning, we suggest adapting models using clinical guidelines and datasets that capture clinicians’ differential diagnosis processes Wang et al. (2025). Future research should also explore efficient fine-tuning strategies that balance performance with computational costs for scenarios requiring frequent model updates in response to evolving medical knowledge.
@@ 7 Conclusion
8 Acknowledgements
We thank the reviewers and the Area Chair for their constructive feedback, which improved this work. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) ProjectID 528483508- FIP 12, as well as the European Union under the grant project 101079894 (COMFORT- Improving Urologic Cancer Care with Artificial Intelligence Solutions), as well as by BMWe SOOFI, Grant Agreement 13IPC040D. The views expressed are solely those of the authors and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HaDEA); neither is responsible for them.
9 Limitations
Our findings should be interpreted within the context of several study-specific limitations. First, our experiments rely on a single large-scale clinical dataset of admission notes, which may not represent the diversity of healthcare systems, patient populations, or note-taking styles. Second, the evaluation focuses on diagnosis prediction from admission-time text and does not assess other clinically relevant outcomes (e.g., treatment recommendations or prognostic reasoning). Third, admission notes frequently exceed the context length of standard encoder models; for encoders with limited attention window, we truncate inputs to the first 512 tokens, which may omit clinically relevant information and bias performance toward evidence appearing early in the document. We partially mitigate this by additionally evaluating a long-context encoder (M2-BERT), but a systematic comparison of long-document strategies (e.g., chunking or hierarchical encoding) is outside the scope of this benchmark. Furthermore, we restrict our evaluation to open-weight models that can be executed locally. Due to the PhysioNet credentialed data use agreement for MIMIC-IV, we do not send protected note text to third-party commercial APIs and therefore do not benchmark closed-source models. As a result, our comparisons do not cover proprietary systems. Finally, while we evaluate multiple architectures and prompting strategies, our hyperparameter choices and model selection may not reflect optimal configurations for every method. Future work should explore broader tuning budgets and additional modeling choices to assess the robustness and generalizability of our conclusions.
10 Ethical Considerations
The use of language models for diagnosis prediction raises important ethical concerns. For example, inaccurate or biased predictions can negatively impact patient care, as admission notes may lack critical information. It is important to note that an automatic diagnosis prediction system should never be seen as a replacement for medical personnel. Rather, it should be used as a supporting tool in the differential diagnosis process. Another major ethical concern is data privacy, as third-party data access can lead to serious consequences, including financial harm, discrimination, and compromised care. Finally, there is a risk that ICD code prediction might be misused to maximize profits, prioritizing financial incentives over patient care. Addressing these concerns is vital for the safe and effective use of AI in healthcare in the future.